OpenStudy (anonymous):

A rectangle has a length that is twice as long as its width and a perimeter of 36. What is its area?

OpenStudy (anonymous):

\[2l+2w=36\] \[2\times 2w+2w=36\] \[6w=36\] \[w=6\]

OpenStudy (anonymous):

area is length times width. width is 6, length is 12

OpenStudy (anonymous):

let the lenght b '2x' n width be 'x' 2(2x+x) = 36 2(3x) = 36 6x = 36 x = 36/6 x = 6 so,width = 6 n lenght = 2*6=12 area = l * b = 12 * 6 = 72

OpenStudy (anonymous):

`2121